8. Determine the mass (in g) of PbCl2 (Ksp, PbCl2 = 1.7x10^-5) that will dissolve in 0.450L of 0.375M Pb(NO3)2 solution.
Answer #1

Given data: Ksp of Pb Cl₂ = 1.78105 Volume of Pb (NO2) = 0.450L Concentration of Pb(NO3)₂ = 0.375m Solution Leat the Pb Cola4.25 8106 = 3 + 0.375 g : > 3 sin omall quantity and Sy is also very small ( so six neglecting 4.25 x156 = 0.375 52 sa = 4.25Solubility of Pb Cl in g/L = Sx molar mass of PbCl2 = 3.366 x 103 malu mal x 270-12 mol = 936-00 X1o 3 g mass of PbCl2 = Solu